# Order-Optimal Global Convergence for Average Reward Reinforcement Learning via Actor-Critic Approach

### v2

@@ 3 PROPOSED ALGORITHM
We propose a Multi-level Monte Carlo-based Natural Actor-Critic (MLMC-NAC) algorithm (Algorithm 1) that runs $K$ number of epochs (also called outer loops). The $k$th loop obtains $\xi_{k}=[\eta_{k},\zeta_{k}]^{\top}$ where $\eta_{k}$ denotes an estimate of the average reward $J(\theta_{k})$, and $\zeta_{k}$ defines an estimate of the critic parameter, $\zeta_{\theta_{k}}$. These estimates are then used to compute the approximate NPG, $\omega_{k}$, which is applied to update the policy parameter $\theta_{k}$.
@@
where $\alpha$ is the learning parameter. The estimate $\xi_{k}$ is obtained in $H$ number of inner loop steps. In particular, $\forall h\in\{0,\cdots,H-1\}$, we apply the following updates starting from arbitrary $\xi_{0}^{k}$, and finally assign $\xi_{H}^{k}=\xi_{k}$.
$\begin{split}&\xi^{k}_{h+1}=\xi^{k}_{h}-\beta\mathbf{v}_{h}(\theta_{k},\xi^{k}_{h}),~{}\text{where}~{}\mathbf{v}_{h}(\theta_{k},\xi^{k}_{h})=\left[c_{\beta}\hat{\nabla}_{\eta}R(\theta_{k},\eta^{k}_{h}),\hat{\nabla}_{\zeta}E(\theta_{k},\xi^{k}_{h})\right]^{\top}\end{split}$ (15)
@@
$\begin{split}\hat{A}^{\pi_{\theta_{k}}}(\xi_{k},z_{t}^{kh})&=r(s_{t}^{kh},a_{t}^{kh})-\eta_{k}+\zeta_{k}^{\top}\left[\phi(s_{t+1}^{kh})-\phi(s_{t}^{kh})\right]\end{split}$ (18)
@@
^$\begin{split}\hat{\nabla}_{\omega}L_{\nu^{\pi_{\theta_{k}}}}(\omega^{k}_{h},\theta_{k},\xi_{k})&=U_{0}+\begin{cases}2^{Q}\left(U^{Q}-U^{Q-1}\right),&\text{if}~{}2^{Q}\leq T_{\max}\\ 0&\text{otherwise}\end{cases}\\ &=\hat{A}^{\mathrm{MLMC}}_{\mathbf{u},k,h}(\theta_{k})\omega_{h}^{k}-\hat{b}^{\mathrm{MLMC}}_{\mathbf{u},k,h}(\theta_{k},\xi_{k})\end{split}$ (20)
@@
$\begin{split}\mathbf{v}^{kh}_{t}&=\begin{bmatrix}c_{\beta}\left\{\eta_{h}^{k}-r(s_{t}^{kh},a_{t}^{kh})\right\}\\ \bigg{\{}(\zeta_{h}^{k})^{\top}\phi(s_{t}^{kh})-\hat{Q}^{\pi_{\theta_{k}}}(\xi_{h}^{k},z_{t}^{kh})\bigg{\}}\phi(s_{t}^{kh})\end{bmatrix}=\hat{A}_{\mathbf{v}}(z_{t}^{kh})\xi^{k}_{h}-\hat{b}_{\mathbf{v}}(z_{t}^{kh})\end{split}$ (21)
@@
$\hat{A}_{\mathbf{v}}(z_{t}^{kh})$ $=\begin{bmatrix}c_{\beta}&0\\ \phi(s_{t}^{kh})&\phi(s_{t}^{kh})\left[\phi(s_{t}^{kh})-\phi(s_{t+1}^{kh})\right]^{\top}\end{bmatrix},$ (22)
@@ 4 MAIN RESULTS
Both Assumptions 2 and 3 are frequently employed in the analysis of actor-critic methods (Suttle et al., 2023; Patel et al., 2024; Wu et al., 2020; Panda and Bhatnagar, 2024). Assumption 2 intuitively relates to the quality of the feature mapping where $\epsilon_{\mathrm{app}}$ measures the quality. Well-designed feature maps may lead to small or even zero $\epsilon_{\mathrm{app}}$, whereas poorly designed features result in higher errors. Assumption 3 is essential for guaranteeing the uniqueness of the solution to the minimization problem (12). Assumption 3 also follows when the set of policy parameters, $\Theta$ is compact and $e\notin W_{\phi}$, where $e$ is the vector of all ones and $W_{\phi}$ is the space spanned by the feature vectors. To see this, note that if $e\notin W_{\phi}$, there exists $\lambda_{\theta}$ for every policy $\pi_{\theta}$ such that $\operatorname{\mathbb{E}}_{\theta}[\phi(s)(\phi(s)-\phi(s^{\prime}))^{\top}]\succcurlyeq\lambda_{\theta}I$ (Zhang et al., 2021b). Since $\Theta$ is compact, setting $\lambda=\inf_{\theta\in\Theta}\lambda_{\theta}>0$ satisfies Assumption 3.
@@
$L_{d^{\pi^{*}}}(\omega_{\theta}^{*};\theta)\leq\epsilon_{\mathrm{bias}}$
@@
$J^{*}-\frac{1}{K}\sum_{k=0}^{K-1}\operatorname{\mathbb{E}}[J(\theta_{k})]$ $\leq\operatorname*{\mathcal{O}}\left(\sqrt{\epsilon_{\mathrm{app}}}+\sqrt{\epsilon_{\mathrm{bias}}}\right)+\tilde{\operatorname*{\mathcal{O}}}\left(t_{\mathrm{mix}}^{3}T^{-1/2}\right)$ (27)
@@ 5 PROOF OUTLINE
We segregate our analysis into four parts as described below.
@@ 5.1 Policy update analysis
$J^{*}-\frac{1}{K}\sum_{k=0}^{K-1}\operatorname{\mathbb{E}}[J(\theta_{k})]\leq\sqrt{\epsilon_{\mathrm{bias}}}$ $+\frac{G_{1}}{K}\sum_{k=0}^{K-1}\operatorname{\mathbb{E}}\|(\operatorname{\mathbb{E}}_{k}\left[\omega_{k}\right]-\omega^{*}_{k})\|+\frac{\alpha G_{2}}{2K}\sum_{k=0}^{K-1}\operatorname{\mathbb{E}}\|\omega_{k}\|^{2}$
$+\frac{1}{\alpha K}\operatorname{\mathbb{E}}_{s\sim d^{\pi^{*}}}[KL(\pi^{*}(\cdot|s)\|\pi_{\theta_{0}}(\cdot|s))]$ (29)
where $KL(\cdot\|\cdot)$ is the Kullback-Leibler divergence, $\omega^{*}_{k}$ is the NPG direction $F(\theta_{k})^{-1}\nabla J(\theta_{k})$, $\pi^{*}$ is the optimal policy, $J^{*}$ is the optimal value of the function $J(\cdot)$, and $\operatorname{\mathbb{E}}_{k}[\cdot]$ denotes conditional expectation given the history up to epoch $k$.
@@
$\begin{split}\operatorname{\mathbb{E}}\|\omega_{k}\|^{2}&\leq 2\operatorname{\mathbb{E}}\|\omega_{k}-\omega_{k}^{*}\|^{2}+2\operatorname{\mathbb{E}}\|\omega_{k}^{*}\|^{2}\overset{(a)}{\leq}2\operatorname{\mathbb{E}}\|\omega_{k}-\omega_{k}^{*}\|^{2}+\dfrac{2}{\mu^{2}}\operatorname{\mathbb{E}}\|\nabla_{\theta}J(\theta_{k})\|^{2}\end{split}$ (30)
@@
$\begin{split}\dfrac{1}{\mu^{2}K}&\left(\sum_{k=0}^{K-1}\|\nabla_{\theta}J(\theta_{k})\|^{2}\right)\leq\dfrac{32LG_{1}^{4}}{\mu^{4}K}+\left(\dfrac{2G_{1}^{4}}{\mu^{2}}+1\right)\left(\dfrac{1}{K}\sum_{k=0}^{K-1}\|\omega_{k}-\omega_{k}^{*}\|^{2}\right)\end{split}$ (31)
@@ 5.2 Analysis of a General Linear Recursion
$\begin{split}&\operatorname{\mathbb{E}}_{h}\left[\left\lVert\hat{P}_{h}-P\right\rVert^{2}\right]\leq\sigma_{P}^{2},~{}\left\lVert\operatorname{\mathbb{E}}_{h}\left[\hat{P}_{h}\right]-P\right\rVert^{2}\leq\delta^{2}_{P},\\ &\operatorname{\mathbb{E}}_{h}\left[\left\lVert\hat{q}_{h}-q\right\rVert^{2}\right]\leq\sigma_{q}^{2},~{}\left\lVert\operatorname{\mathbb{E}}_{h}\left[\hat{q}_{h}\right]-q\right\rVert^{2}\leq\delta_{q}^{2},~{}\text{and~{}}\left\lVert\operatorname{\mathbb{E}}\left[\hat{q}_{h}\right]-q\right\rVert^{2}\leq\bar{\delta}_{q}^{2}\end{split}$ (33)
@@
$\|\operatorname{\mathbb{E}}[x_{H}]-x^{*}\|^{2}\leq\frac{\|\operatorname{\mathbb{E}}[x_{0}]-x^{*}\|^{2}}{H^{2}}+\operatorname*{\mathcal{O}}(\bar{R}_{1})+\operatorname*{\mathcal{O}}\left(\lambda_{P}^{-2}\delta_{P}^{2}\left\{\operatorname{\mathbb{E}}\left[\|x_{0}-x^{*}\|^{2}\right]+\mathcal{O}\left(R_{0}+R_{1}\right)\right\}\right)$
@@ 5.3 Analysis of NPG-Finding Subroutine
$(b)~{}\operatorname{\mathbb{E}}_{k,h}\left[\left\lVert\hat{A}^{\mathrm{MLMC}}_{\mathbf{u},k,h}(\theta_{k})-F(\theta_{k})\right\rVert^{2}\right]\leq\operatorname*{\mathcal{O}}\left(G_{1}^{4}t_{\mathrm{mix}}\log T_{\max}\right)$
$(c)~{}\left\lVert\operatorname{\mathbb{E}}_{k,h}\left[\hat{b}^{\mathrm{MLMC}}_{\mathbf{u},k,h}(\theta_{k},\xi_{k})\right]-\nabla_{\theta}J(\theta_{k})\right\rVert^{2}\leq\operatorname*{\mathcal{O}}\left(\sigma_{\mathbf{u},k}^{2}t_{\mathrm{mix}}T_{\max}^{-1}+\delta_{\mathbf{u},k}^{2}\right)$
$(d)~{}\operatorname{\mathbb{E}}_{k,h}\left[\left\lVert\hat{b}^{\mathrm{MLMC}}_{\mathbf{u},k,h}(\theta_{k},\xi_{k})-\nabla_{\theta}J(\theta_{k})\right\rVert^{2}\right]\leq\operatorname*{\mathcal{O}}\left(\sigma_{\mathbf{u},k}^{2}t_{\mathrm{mix}}\log T_{\max}+\delta_{\mathbf{u},k}^{2}\right)$
@@
$\begin{split}&(e)~{}\left\lVert\operatorname{\mathbb{E}}_{k}\left[\hat{b}^{\mathrm{MLMC}}_{\mathbf{u},k,h}(\theta_{k},\xi_{k})\right]-\nabla_{\theta}J(\theta_{k})\right\rVert^{2}\leq\operatorname*{\mathcal{O}}\left(\bar{\sigma}_{\mathbf{u},k}^{2}t_{\mathrm{mix}}T_{\max}^{-1}+\bar{\delta}_{\mathbf{u},k}^{2}\right)\end{split}$
@@
$\mathbb{E}_{k}\left[\|\omega_{k}-\omega_{k}^{*}\|^{2}\right]\leq\frac{1}{H^{2}}\|\omega^{k}_{H}-\omega_{k}^{*}\|^{2}+\tilde{\operatorname*{\mathcal{O}}}\left(\frac{G_{1}^{6}t_{\mathrm{mix}}^{3}}{\mu^{4}H}+\dfrac{G_{1}^{2}c_{\beta}^{2}t_{\mathrm{mix}}}{\mu^{2}\lambda^{2}H}\right)+\mu^{-2}G_{1}^{2}\operatorname*{\mathcal{O}}\big{(}\operatorname{\mathbb{E}}_{k}\|\xi_{k}-\xi_{k}^{*}\|^{2}+\epsilon_{\mathrm{app}}\big{)}$
@@
$\|\mathbb{E}_{k}[\omega_{k}]-\omega_{k}^{*}\|^{2}$ $\leq\tilde{\operatorname*{\mathcal{O}}}\left(\dfrac{G_{1}^{4}t_{\mathrm{mix}}}{\mu^{2}H^{2}}\left\lVert\omega_{H}^{k}-\omega_{k}^{*}\right\rVert^{2}\right)+\mu^{-2}G_{1}^{2}\operatorname*{\mathcal{O}}\bigg{(}\|\operatorname{\mathbb{E}}_{k}[\xi_{k}]-\xi_{k}^{*}\|^{2}+\epsilon_{\mathrm{app}}\bigg{)}$
$+\tilde{\operatorname*{\mathcal{O}}}\left(\dfrac{G_{1}^{6}t^{2}_{\mathrm{mix}}}{\mu^{4}H^{2}}\operatorname{\mathbb{E}}_{k}\left[\left\lVert\xi_{k}-\xi_{k}^{*}\right\rVert^{2}\right]\right)+\tilde{\operatorname*{\mathcal{O}}}\left(\dfrac{G_{1}^{4}t_{\mathrm{mix}}}{\mu^{2}H^{2}}\left\{\mu^{-4}G_{1}^{6}t_{\mathrm{mix}}^{3}+\mu^{-2}\lambda^{-2}G_{1}^{2}c_{\beta}^{2}t_{\mathrm{mix}}\right\}\right)$

### v3

@@ 3 Proposed Algorithm
We propose a Multi-level Monte Carlo-based Natural Actor-Critic (MLMC-NAC) algorithm (Algorithm 1) that runs $K$ number of epochs (also called outer loops). The $k$th loop obtains $\xi_{k}=[\eta_{k},\zeta_{k}]^{\top}$ where $\eta_{k}$ denotes an estimate of the average reward $J(\theta_{k})$, and $\zeta_{k}$ is an estimate of the critic parameter, $\zeta_{\theta_{k}}$. These estimates are then used to compute the approximate NPG, $\omega_{k}$, which is applied to update the policy parameter $\theta_{k}$.
@@
where $\alpha$ is a learning parameter. The estimate $\xi_{k}$ is obtained in $H$ inner loop steps. In particular, $\forall h\in\{0,\cdots,H-1\}$, we apply the following updates starting from arbitrary $\xi_{0}^{k}$, and finally assign $\xi_{H}^{k}=\xi_{k}$.
$\begin{split}&\xi^{k}_{h+1}=\xi^{k}_{h}-\beta\mathbf{v}_{h}(\theta_{k},\xi^{k}_{h}),~{}\text{where}\\ ~{}&\mathbf{v}_{h}(\theta_{k},\xi^{k}_{h})=\left[c_{\beta}\hat{\nabla}_{\eta}R(\theta_{k},\eta^{k}_{h}),\hat{\nabla}_{\zeta}E(\theta_{k},\xi^{k}_{h})\right]^{\top}\end{split}$ (16)
@@
$\begin{split}\hat{A}^{\pi_{\theta_{k}}}(\xi_{k},z_{t}^{kh})&=r(s_{t}^{kh},a_{t}^{kh})-\eta_{k}\\ &\hskip 28.45274pt+\zeta_{k}^{\top}\left[\phi(s_{t+1}^{kh})-\phi(s_{t}^{kh})\right]\end{split}$ (19)
@@
^$\begin{split}\hat{\nabla}_{\omega}&L_{\nu^{\pi_{\theta_{k}}}}(\omega^{k}_{h},\theta_{k},\xi_{k})\\ &=U_{0}+\begin{cases}2^{Q}\left(U^{Q}-U^{Q-1}\right),&\text{if}~{}2^{Q}\leq T_{\max}\\ 0&\text{otherwise}\end{cases}\\ &=\hat{A}^{\mathrm{MLMC}}_{\mathbf{u},k,h}(\theta_{k})\omega_{h}^{k}-\hat{b}^{\mathrm{MLMC}}_{\mathbf{u},k,h}(\theta_{k},\xi_{k})\end{split}$ (21)
@@
$\begin{split}\mathbf{v}^{kh}_{t}&=\begin{bmatrix}c_{\beta}\left\{\eta_{h}^{k}-r(s_{t}^{kh},a_{t}^{kh})\right\}\\ \bigg{\{}(\zeta_{h}^{k})^{\top}\phi(s_{t}^{kh})-\hat{Q}^{\pi_{\theta_{k}}}(\xi_{h}^{k},z_{t}^{kh})\bigg{\}}\phi(s_{t}^{kh})\end{bmatrix}\\ &=\hat{A}_{\mathbf{v}}(z_{t}^{kh})\xi^{k}_{h}-\hat{b}_{\mathbf{v}}(z_{t}^{kh})\end{split}$ (22)
@@
[…]
@@ 4 Main Results
Both Assumptions 2 and 3 are frequently employed in the analysis of actor-critic methods (Suttle et al., 2023; Patel et al., 2024; Wu et al., 2020; Panda & Bhatnagar, 2024). Assumption 2 intuitively relates to the quality of the feature mapping where $\epsilon_{\mathrm{app}}$ measures the quality. Well-designed feature maps may lead to small or even zero $\epsilon_{\mathrm{app}}$, whereas poorly designed features result in higher errors. Assumption 3 is essential for guaranteeing the uniqueness of the solution to the minimization problem (13). Assumption 3 also follows when the set of policy parameters, $\Theta$ is compact and $e\notin W_{\phi}$, where $e$ is the vector of all ones and $W_{\phi}$ is the space spanned by the feature vectors. To see this, note that if $e\notin W_{\phi}$, there exists $\lambda_{\theta}$ for every policy $\pi_{\theta}$ such that $\operatorname{\mathbb{E}}_{\theta}[\phi(s)(\phi(s)-\phi(s^{\prime}))^{\top}]\succcurlyeq\lambda_{\theta}I$ (Zhang et al., 2021b). Since $\Theta$ is compact, setting $\lambda=\inf_{\theta\in\Theta}\lambda_{\theta}>0$ satisfies Assumption 3.
@@
$L_{\nu^{\pi^{*}}}(\omega_{\theta}^{*};\theta)\leq\epsilon_{\mathrm{bias}}$
@@
$J^{*}-\frac{1}{K}\sum_{k=0}^{K-1}\operatorname{\mathbb{E}}[J(\theta_{k})]$ $\leq\operatorname*{\mathcal{O}}\left(\sqrt{\epsilon_{\mathrm{app}}}+\sqrt{\epsilon_{\mathrm{bias}}}\right)$
$\hskip 28.45274pt+\tilde{\operatorname*{\mathcal{O}}}\left(t_{\mathrm{mix}}^{3}T^{-1/2}\right)$ (28)
@@ 5.1 Policy update analysis
$J^{*}-\frac{1}{K}\sum_{k=0}^{K-1}\operatorname{\mathbb{E}}[J(\theta_{k})]\leq\sqrt{\epsilon_{\mathrm{bias}}}$
$+\frac{G_{1}}{K}\sum_{k=0}^{K-1}\operatorname{\mathbb{E}}\|(\operatorname{\mathbb{E}}_{k}\left[\omega_{k}\right]-\omega^{*}_{k})\|+\frac{\alpha G_{2}}{2K}\sum_{k=0}^{K-1}\operatorname{\mathbb{E}}\|\omega_{k}\|^{2}$
$+\frac{1}{\alpha K}\operatorname{\mathbb{E}}_{s\sim d^{\pi^{*}}}[\mathrm{KL}(\pi^{*}(\cdot|s)\|\pi_{\theta_{0}}(\cdot|s))]$ (30)
where $\mathrm{KL}(\cdot\|\cdot)$ is the Kullback-Leibler divergence, $\omega^{*}_{k}$ is the NPG direction $F(\theta_{k})^{-1}\nabla J(\theta_{k})$, $\pi^{*}$ is the optimal policy, $J^{*}$ is the optimal value of the function $J(\cdot)$, and $\operatorname{\mathbb{E}}_{k}[\cdot]$ denotes conditional expectation given the history up to epoch $k$.
@@
$\begin{split}\operatorname{\mathbb{E}}\|\omega_{k}\|^{2}&\leq 2\operatorname{\mathbb{E}}\|\omega_{k}-\omega_{k}^{*}\|^{2}+2\operatorname{\mathbb{E}}\|\omega_{k}^{*}\|^{2}\\ &\overset{(a)}{\leq}2\operatorname{\mathbb{E}}\|\omega_{k}-\omega_{k}^{*}\|^{2}+\dfrac{2}{\mu^{2}}\operatorname{\mathbb{E}}\|\nabla_{\theta}J(\theta_{k})\|^{2}\end{split}$ (31)
@@
$\begin{split}\dfrac{1}{\mu^{2}K}&\left(\sum_{k=0}^{K-1}\|\nabla_{\theta}J(\theta_{k})\|^{2}\right)\leq\dfrac{32LG_{1}^{4}}{\mu^{4}K}\\ &+\left(\dfrac{2G_{1}^{4}}{\mu^{2}}+1\right)\left(\dfrac{1}{K}\sum_{k=0}^{K-1}\|\omega_{k}-\omega_{k}^{*}\|^{2}\right)\end{split}$ (32)
@@ 5.2 Analysis of a General Linear Recursion
$\begin{split}&\operatorname{\mathbb{E}}_{h}\left[\left\lVert\hat{P}_{h}-P\right\rVert^{2}\right]\leq\sigma_{P}^{2},~{}\left\lVert\operatorname{\mathbb{E}}_{h}\left[\hat{P}_{h}\right]-P\right\rVert^{2}\leq\delta^{2}_{P},\\ &\operatorname{\mathbb{E}}_{h}\left[\left\lVert\hat{q}_{h}-q\right\rVert^{2}\right]\leq\sigma_{q}^{2},~{}\left\lVert\operatorname{\mathbb{E}}_{h}\left[\hat{q}_{h}\right]-q\right\rVert^{2}\leq\delta_{q}^{2},\\ &\text{and~{}}\left\lVert\operatorname{\mathbb{E}}\left[\hat{q}_{h}\right]-q\right\rVert^{2}\leq\bar{\delta}_{q}^{2}\end{split}$ (34)
@@
$\|\operatorname{\mathbb{E}}[x_{H}]-x^{*}\|^{2}\leq\frac{\|\operatorname{\mathbb{E}}[x_{0}]-x^{*}\|^{2}}{H^{2}}+\operatorname*{\mathcal{O}}(\bar{R}_{1})$
$+\operatorname*{\mathcal{O}}\left(\lambda_{P}^{-2}\delta_{P}^{2}\left\{\operatorname{\mathbb{E}}\left[\|x_{0}-x^{*}\|^{2}\right]+\mathcal{O}\left(R_{0}+R_{1}\right)\right\}\right)$
@@ 5.3 Analysis of NPG-Finding Subroutine
$(b)~{}\operatorname{\mathbb{E}}_{k,h}\left[\left\lVert\hat{A}^{\mathrm{MLMC}}_{\mathbf{u},k,h}(\theta_{k})-F(\theta_{k})\right\rVert^{2}\right]$
$\hskip 85.35826pt\leq\operatorname*{\mathcal{O}}\left(G_{1}^{4}t_{\mathrm{mix}}\log T_{\max}\right)$
$(c)~{}\left\lVert\operatorname{\mathbb{E}}_{k,h}\left[\hat{b}^{\mathrm{MLMC}}_{\mathbf{u},k,h}(\theta_{k},\xi_{k})\right]-\nabla_{\theta}J(\theta_{k})\right\rVert^{2}$
$\hskip 85.35826pt\leq\operatorname*{\mathcal{O}}\left(\sigma_{\mathbf{u},k}^{2}t_{\mathrm{mix}}T_{\max}^{-1}+\delta_{\mathbf{u},k}^{2}\right)$
$(d)~{}\operatorname{\mathbb{E}}_{k,h}\left[\left\lVert\hat{b}^{\mathrm{MLMC}}_{\mathbf{u},k,h}(\theta_{k},\xi_{k})-\nabla_{\theta}J(\theta_{k})\right\rVert^{2}\right]$
$\hskip 85.35826pt\leq\operatorname*{\mathcal{O}}\left(\sigma_{\mathbf{u},k}^{2}t_{\mathrm{mix}}\log T_{\max}+\delta_{\mathbf{u},k}^{2}\right)$
@@
$\begin{split}&(e)~{}\left\lVert\operatorname{\mathbb{E}}_{k}\left[\hat{b}^{\mathrm{MLMC}}_{\mathbf{u},k,h}(\theta_{k},\xi_{k})\right]-\nabla_{\theta}J(\theta_{k})\right\rVert^{2}\\ &\hskip 85.35826pt\leq\operatorname*{\mathcal{O}}\left(\bar{\sigma}_{\mathbf{u},k}^{2}t_{\mathrm{mix}}T_{\max}^{-1}+\bar{\delta}_{\mathbf{u},k}^{2}\right)\end{split}$
@@
$\mathbb{E}_{k}\left[\|\omega_{k}-\omega_{k}^{*}\|^{2}\right]\leq\frac{1}{H^{2}}\|\omega^{k}_{H}-\omega_{k}^{*}\|^{2}+\tilde{\operatorname*{\mathcal{O}}}\bigg{(}\frac{G_{1}^{6}t_{\mathrm{mix}}^{3}}{\mu^{4}H}\bigg{)}$
$\tilde{\operatorname*{\mathcal{O}}}\left(\dfrac{G_{1}^{2}c_{\beta}^{2}t_{\mathrm{mix}}}{\mu^{2}\lambda^{2}H}\right)+\mu^{-2}G_{1}^{2}\operatorname*{\mathcal{O}}\big{(}\operatorname{\mathbb{E}}_{k}\|\xi_{k}-\xi_{k}^{*}\|^{2}+\epsilon_{\mathrm{app}}\big{)}$
@@
$\|\mathbb{E}_{k}[\omega_{k}]-\omega_{k}^{*}\|^{2}\leq\tilde{\operatorname*{\mathcal{O}}}\left(\dfrac{G_{1}^{4}t_{\mathrm{mix}}}{\mu^{2}H^{2}}\left\lVert\omega_{H}^{k}-\omega_{k}^{*}\right\rVert^{2}\right)$
$+\mu^{-2}G_{1}^{2}\operatorname*{\mathcal{O}}\bigg{(}\|\operatorname{\mathbb{E}}_{k}[\xi_{k}]-\xi_{k}^{*}\|^{2}+\epsilon_{\mathrm{app}}\bigg{)}$
$+\tilde{\operatorname*{\mathcal{O}}}\left(\dfrac{G_{1}^{6}t^{2}_{\mathrm{mix}}}{\mu^{4}H^{2}}\operatorname{\mathbb{E}}_{k}\left[\left\lVert\xi_{k}-\xi_{k}^{*}\right\rVert^{2}\right]\right)$
$+\tilde{\operatorname*{\mathcal{O}}}\left(\dfrac{G_{1}^{4}t_{\mathrm{mix}}}{\mu^{2}H^{2}}\left\{\mu^{-4}G_{1}^{6}t_{\mathrm{mix}}^{3}+\mu^{-2}\lambda^{-2}G_{1}^{2}c_{\beta}^{2}t_{\mathrm{mix}}\right\}\right)$